What is the boiling point of a 1.5 solution of MgSO4 in water ( Kf=1.86 C, Kb= 0.512


1
Expert's answer
2021-02-04T07:07:20-0500

Ulike UNIUNT UT lotur ophoniane blube - homo 2) How much MgSO4 (in g) would you dissolve in 100 g water to prepare 1.5 molality MgSO4(aq) solution? (Show your work to get full credit) Mass of solvet 1000 100K Cork ndanits = number of moks 1000 Mass of solvent Molaris=1.5m Molar mass of my so = 24.39+ 329 + 98 Cheg) = 120. 39 mass of masos - 1.5x 1203 xoil To. 0459 3) Calculate the freezing and boiling point for the solution above. [k (H2O) = 1.86 °C/m, kb(H2O) = 0.512 °C/m] (Show your work to get full credit) HEY 4) Under which of the following sets of conditions could the most O2(g) be dissolved in H2O(l)? a. high pressure above the solution and low temperature The higher me teenopre presun b. low pressure above the solution and low temperature Touer temp c. high pressure above the solution and high temperature d low nressure above the solution and low temperature untur does not affect was solubility